# Hetero-EUCLID: Interpretable model discovery for heterogeneous hyperelastic materials using stress-unsupervised learning11footnotemark: 1

### v1

@@ Abstract
We propose a computational framework, Hetero-EUCLID, for segmentation and parameter identification to characterize the full hyperelastic behavior of all constituents of a heterogeneous material. In this work, we leverage the Bayesian-EUCLID (Efficient Unsupervised Constitutive Law Identification and Discovery) framework to efficiently solve the heterogenized formulation through parsimonious model selection using sparsity-promoting priors and Monte Carlo Markov Chain sampling. We utilize experimentally observable 3D surface displacement and boundary-averaged force data generated from Finite Element simulations of non-equi-biaxial tension tests on heterogeneous specimens. The framework broadly consists of two steps– residual force-based segmentation, and constitutive parameter identification. We validate and demonstrate the ability of the proposed framework to segment the domain, and characterize the constituent materials on various types of thin square heterogeneous domains. We validate of the framework’s ability to segment and characterize materials with various levels of displacement noises and non-native mesh discretizations, i.e, using different meshes for the forward FE simulations and the inverse EUCLID problem. This demonstrates Hetero-EUCLID framework’s applicability in Digital Image/Volume Correlation-based experimental scenarios. Furthermore, the proposed framework performs successful segmentation and material characterizations based on data from a single experiment, thereby making it viable for rapid, interpretable model discovery in domains such as aerospace and defense composites and for characterization of selective tissue stiffening in medical conditions such as fibroatheroma, atherosclerosis, or cancer.
@@ 1 Introduction
With the increasing relevance of biomedical engineering, there is a growing focus on understanding the mechanical behavior of biological tissues. Many of these tissues exhibit intrinsic material heterogeneity [1, 2, 3], necessitating the development of multi-segmented constitutive models. Such models can significantly enhance our understanding of complex biomechanical processes, including the progression of aortic aneurysms [4, 5, 6] and the localized stiffening of tissues due to tumor growth [7, 8].
@@
A more traditional approach to constitutive modeling had been to perform the phenomenological model calibration of an apriori assumed constitutive equation, utilizing the full-field displacement data obtained from Digital Image Correlation (DIC) in combination with boundary-averaged force measurement from force transducers [9]. The inverse problem of determining the constitutive parameters is solved using Finite Element Model Updating (FEMU) [10] or Virtual Fields Method (VFM) [11]. However, assuming a form of the constitutive equation for the material limits the materials and constitutive models that can be characterized with this approach. To address the limitations of traditional model calibration, and enable characterization and discovery of materials that are not necessarily well explained by existing models, several works in the past have sought to entirely bypass [12, 13, 14, 15, 16, 17, 18, 19] or surrogate constitutive models. The approach of surrogating constitutive model identifies a relationship between mechanical stresses and strains using Gaussian Process Regression (GPR) [20, 21], Radial Basis Functions (RBF) [22], piece-wise polynomial interpolation [23, 24], and Artificial Neural Networks (ANN) [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. Furthermore, to address the lack of availability of full stress-tensor data for training supervised training several frameworks leverage the balance of linear momentum [36] or the virtual fields method [37] to train ANNs or Recurrent Neural Networks (RNNs) to surrogate the stress-strain relationship for the materials using only experimentally available surface displacement and boundary-averaged force data.
@@
Although stress-unsupervised ANN/RNN surrogates are capable of learning complicated constitutive models, they do not provide interpretable constitutive models. The EUCLID framework [38] was initially introduced to model complicated homogeneous hyperelastic constitutive behavior in an interpretable manner. This was achieved through sparse symbolic regression using a pre-defined library of known strain energy densities. The EUCLID framework has since been extended to model homogeneous elastoplastic materials [39], viscoelastic materials [40], and generalized standard materials [41]. Most of the aforementioned works address the modeling of homogeneous materials.
@@
An earlier work which used only the surface displacement data [42] focused on elastography of incompressible linear elastic heterogeneous solids using novel Physics Informed Neural Networks (PINNs) to predict the variation of the Young’s Modulus as a function of position in the material. A later work by the same authors employed multiple PINNs to obtain the spatial variation of Poisson’s ratio in elastography [43]. Another work presented a detailed study on segmenting heterogeneous hyperelastic materials using surface displacement data and associated strain invariants [44]. A more recent work used a stress-unsupervised ANN-based framework [45] to segment and model a heterogeneous hyperelastic solid, wherein the inputs to the ANN are the strain invariants and the query position, and the output of the ANN is the strain energy density. The novel ANN-based framework was able to segment and reproduce the constitutive models of each segment in the hyperelastic solid. However, the use of ANNs makes the discovered models uninterpretable and provides no uncertainty estimates on the discovered models. Furthermore, since the query position is an essential input to the aforementioned ANN, it would complicate the ANN’s adoption as a constitutive model for forward simulation on other geometries. In this context, we propose the Hetero-EUCLID framework, wherein we extend the Bayesian-EUCLID framework [46] to segment a heterogeneous solid and provide interpretable constitutive models for all constituents, using only surface displacement and boundary-averaged force data.
@@
The problem of constitutively modeling all the constituents in a heterogeneous geometry can be divided into two subtasks: segmenting the geometry and obtaining the constitutive model for all the materials involved. In section 2 of the manuscript we present
@@
  • An overview of extending the Bayesian-EUCLID framework [46], which was originally developed for homogeneous materials, to model segmented homogeneous materials.
@@
A schematic illustration of the proposed Hetero-EUCLID framework is provided in Figure 1. In section 3, we present results of validation studies on square geometries containing two or more hyperelastic constituents. The validation studies consist of using surface displacement data and boundary force data from biaxial tension tests on thin hyperelastic specimens. For the scope of this study, we source displacement and boundary-averaged force data from 3D Finite Element Simulations. This is in contrast to previous similar studies, which focused on validating their results on 2D plane strain settings [36, 45]. In section 3 we also validate the proposed Hetero-EUCLID framework on “Non-native” meshed geometries, wherein the Finite Element mesh used to perform the inverse Hetero-EUCLID analysis was different from the Finite Element mesh used to generate the displacement and boundary-averaged force. The manuscript concludes in section 4 with comments on the results of the study and potential future work.
Figure 1: Schematic of the Hetero-EUCLID workflow for unsupervised constitutive model discovery in heterogeneous hyperelastic materials. (a) illustrates the specimen subjected to non-equi-biaxial tensile loading. (b) and (c) show the point-wise displacement observations and corresponding finite element interpolation onto a wedge-type mesh, respectively. (d) depicts the resulting full-field displacement field, while (e) presents the corresponding strain field obtained via spatial differentiation. (f) represents a catalogue of the hyperelastic constitutive models library. (g) outlines the homogenized model formulation incorporating the weak form of the linear momentum balance, which is solved using ordinary least squares regression to yield the residual force norm distribution shown in (h). Material segmentation based on the residual force indicator is performed using a random seed-based island growth algorithm, as illustrated in (i). (j) presents the unified global heterogenized system of equations assembled after material segmentation. The unknown constitutive models $\boldsymbol{\theta}^{(i)}$ are solved for using the Bayesian EUCLID framework. (k) highlights the discovered region-wise interpretable constitutive models corresponding to each material segment within the heterogeneous specimen.
@@ 2.1 Incorporating heterogeneity
Figure 2: Meshing the given thin plate geometry with C3D6 elements that have a thickness identical to that of the plate.
@@
where $\nabla(\cdot)=\partial(\cdot)/\partial\boldsymbol{X}$ is defined as the gradient operator in the reference coordinates. Due to the use of linear shape functions in the wedge mesh type, $\nabla\mathbb{N}^{a}(X)$ remains constant within each element. For the scope of this work, a single-element-thick mesh is generated over the reference domain. This would imply that all nodes are located on the surface and their displacements $\boldsymbol{u}^{a,t}$ could be obtained through DIC experiments. Further, for brevity, in this study we utilize data from a single large deformation, i.e, $t=1$. We will henceforth be dropping the use of $t$ in the superscript.
@@
We now consider a reference domain composed of multiple hyperelastic material segments. Since each segment belongs to the hyperelastic material class, they can be assigned a strain energy density $W(\boldsymbol{F})$ and the corresponding first Piola–Kirchhoff stress tensor would be $\boldsymbol{P}(\boldsymbol{F})=\partial W/\partial\boldsymbol{F}$. In order to satisfy the material objectivity constraint, the strain energy density will be a function of the right Cauchy-Green deformation tensor $\boldsymbol{C}=\boldsymbol{F}^{T}\boldsymbol{F}$. Furthermore, given that we focus on isotropic materials for this study, the strain energy density $W(\boldsymbol{C})$ can be defined in terms of isotropic invariants $(I_{1},I_{2},I_{3})$ as
$W(\boldsymbol{C})=W(I_{1},I_{2},I_{3})$ (3)
where,
$I_{1}(\boldsymbol{C})=\mathrm{tr}(\boldsymbol{C}),\quad I_{2}(\boldsymbol{C})=\frac{1}{2}[\mathrm{tr}^{2}(\boldsymbol{C})-\mathrm{tr}(\boldsymbol{C}^{2})],\quad I_{3}(\boldsymbol{C})=\mathrm{det}(\boldsymbol{C})$ (4)
@@
We address the inverse problem of identifying the strain energy density function for a specimen subjected to a non-equi-biaxial loading test by introducing a material feature library $\boldsymbol{Q}:\mathbb{R}^{3}\rightarrow\mathbb{R}^{\mathit{n_{f}}}$, where $\mathit{n_{f}}$ denotes the number of nonlinear isotropic feature functions. Accordingly, the constitutive relation in Equation 3 can be reformulated using this feature library $\boldsymbol{Q}$ and a vector of unknown material parameters $\boldsymbol{\theta}\in\mathbb{R}^{\mathit{n_{f}}}_{+}$, consisting of non-negative scalar coefficients. Following [46], we impose the constraint on each component of $\boldsymbol{\theta}$ to be non-negative to ensure polyconvexity of $W(I_{1},I_{2},I_{3})$. It should be noted that any terms involving $(\tilde{I}_{2}-3)$ in the feature library $\boldsymbol{Q}$ are not polyconvex, and should ideally be replaced with $(\tilde{I}_{2}^{3/2}-3^{3/2})$ [47]. However, we choose to retain the more traditional $(\tilde{I}_{2}-3)$ term, which appears in the Isihara model [48], and is not known to cause any physical admissibility issues.
$W(I_{1},I_{2},I_{3};\boldsymbol{\theta})=\boldsymbol{Q}^{T}(I_{1},I_{2},I_{3})\>\boldsymbol{\theta}$ (5)
@@
For the scope of this work to demonstrate the capability of identifying the heterogeneity in the specimen, we have appropriately chosen the following feature library, which can represent the compressible hyperelastic isotropic material models such as generalized Mooney-Rivlin, which utilizes the isotropic invariants ($I_{1},I_{2},I_{3}$) covering Neo-Hookean [49], Isihara [48], Biderman [50], and Haines-Wilson [51].
@@
$\boldsymbol{Q}\left(I_{1},I_{2},I_{3}\right)$ $=\underbrace{\left[\left(\tilde{I}_{1}-3\right)^{i}\left(\tilde{I}_{2}-3\right)^{j-i}:j\in\{1,\dots,N\},i\in\{0,\dots,j\}\right]^{T}}_{\text{Generalized Mooney-Rivlin features }}$
@@
$\quad\oplus\quad\underbrace{\left[(J-1)^{2k}:k\in\{1,\dots,M\}\right]^{T}}_{\text{Volumetric energy features }}$ (6)
where $\tilde{I}_{1}=J^{-2/3}I_{1},\>\tilde{I}_{2}=J^{-4/3}I_{2},\>J=\mathrm{det}(\boldsymbol{F})=I_{3}^{1/2}$ are the isotropic invariants of the Cauchy-Green deformation tensor $(\boldsymbol{C})$ and the symbol $\oplus$ represents the concatenation of the features of the chosen library. It is important to note that the requirements of satisfying physical constraints such as objectivity, material symmetry, and stress-free reference configuration ($\boldsymbol{P}(\boldsymbol{F}=\boldsymbol{I})=0$) are automatically satisfied by Equation 5 and 6.
@@
To compute the partial derivatives of the various features in the material library, $\partial\boldsymbol{Q}^{T}/\partial F_{i\mathrm{J}}$, we use the automatic differentiation [52] feature in the torch library of Python. An advantage of using the weak form of the linear momentum balance is that it inherently reduces the sensitivity issues associated with the second-order spatial derivatives present in the strong form [53].
@@
Now, we choose various nodal values of the test function $\boldsymbol{v}^{a}$ to generate equations to solve for the constitutive properties of the material.
@@
    Within each integral, the material model $\boldsymbol{\theta}\left(\boldsymbol{X}\right)=\boldsymbol{\theta^{(k)}},\text{for }\boldsymbol{X}\in\Omega^{(k)};k\in\{1,\dots,n_{c}\}$. Since $\boldsymbol{\theta^{(k)}}$ is constant in each material segment it can be taken out of the integral. Equation 13 can thus be rewritten as
@@
    $v_{i}^{a}=\begin{cases}1&\forall\>(a,i)\in\mathcal{D}^{\mathrm{fix}}_{k}\\ 0&\text{otherwise}\end{cases}\quad\text{for }k=\{1,\dots,n_{b}\}$
    where, $\mathcal{D}_{k}^{\text{fix}}\subset\mathcal{D}^{\text{fix}}$ denotes the set of degree of freedoms belonging to the $k^{\text{th}}$ boundary $\partial\Omega_{k}$ and $n_{b}$ denotes the total number of boundaries where force is measured. Substituting the value of chosen test function $v_{i}^{a}$, Equation 10 gets modified as follows
@@
    Usually, in experiments, load cells measure only one component of force, i.e, normal force. In this case, only component $i\in\{1,2,3\}$ of $R_{i}^{k}$ would be recorded at a boundary $k$. However, specialized shear load cells can also be used to measure the other two components. In this study, for convenience, we assume all three components of $\boldsymbol{R}^{k}$ are known at every boundary $k\in\{1\dots n_{b}\}$. Thus, using Equation 8 and the methodology adopted in Equations 13-15, Equation 17 can be written as

### v2

# Heterogeneous material , data-driven discovery , hyperelastic constitutive models , Bayesian learning , uncertainty quantification journal: arXiv

@@ Abstract
We propose a computational framework, Hetero-EUCLID, for segmentation and parameter identification to characterize the full hyperelastic behavior of all constituents of a heterogeneous material. In this work, we leverage the Bayesian-EUCLID (Efficient Unsupervised Constitutive Law Identification and Discovery) framework to efficiently solve the heterogenized formulation through parsimonious model selection using sparsity-promoting priors and Monte Carlo Markov Chain sampling. We utilize experimentally observable 3D surface displacement and boundary-averaged force data generated from Finite Element simulations of non-equi-biaxial tension tests on heterogeneous specimens. The framework broadly consists of two steps– residual force-based segmentation and constitutive parameter identification. We validate and demonstrate the ability of the proposed framework to segment the domain and characterize the constituent materials on various types of thin square heterogeneous domains. We validate the framework’s ability to segment and characterize materials with multiple levels of displacement noises and non-native mesh discretizations, i.e, using different meshes for the forward FE simulations and the inverse EUCLID problem. This demonstrates the applicability of the Hetero-EUCLID framework in Digital Image/Volume Correlation-based experimental scenarios. Furthermore, the proposed framework performs successful segmentation and material characterizations based on data from a single experiment, thereby making it viable for rapid, interpretable model discovery in domains such as aerospace and defense composites and for characterization of selective tissue stiffening in medical conditions such as fibroatheroma, atherosclerosis, or cancer.
@@ 1 Introduction
With the increasing relevance of biomedical engineering, there is a growing focus on understanding the mechanical behavior of biological tissues. Many of these tissues exhibit intrinsic material heterogeneity [Guo et al., 2022, Wei et al., 2023, Lee et al., 2024], necessitating the development of multi-segmented constitutive models. Such models can significantly enhance our understanding of complex biomechanical processes, including the progression of aortic aneurysms [Burris et al., 2022, Lin et al., 2022, Vander Linden et al., 2023] and the localized stiffening of tissues due to tumor growth [Goriely, 2017, Micalet et al., 2021].
@@
A more traditional approach to constitutive modeling involves performing phenomenological model calibration of an a priori assumed constitutive equation, utilizing full-field displacement data obtained from Digital Image Correlation (DIC) in combination with boundary-averaged force measurements from force transducers [Hild and Roux, 2006]. The inverse problem of determining the constitutive parameters is solved using Finite Element Model Updating (FEMU) [Marwala, 2010] or Virtual Fields Method (VFM) [Grediac et al., 2006]. However, assuming a form of the constitutive equation for the material limits the materials and constitutive models that can be characterized with this approach. To address the limitations of traditional model calibration, and enable characterization and discovery of materials that are not necessarily well explained by existing models, several works in the past have sought to entirely bypass [Kirchdoerfer and Ortiz, 2016, Ibañez et al., 2017, Kirchdoerfer and Ortiz, 2018, Conti et al., 2018, Nguyen and Keip, 2018, Eggersmann et al., 2019, Carrara et al., 2020, Karapiperis et al., 2021], or surrogate constitutive models. The approach of surrogating constitutive model identifies a relationship between mechanical stresses and strains using Gaussian Process Regression (GPR) [Rocha et al., 2020, Fuhg et al., 2022], Radial Basis Functions (RBF) [Park et al., 2021], piece-wise polynomial interpolation [Crespo et al., 2017, Sussman and Bathe, 2009], and Artificial Neural Networks (ANN) [Ghaboussi et al., 1991, Fernández et al., 2021, Klein et al., 2022, Vlassis and Sun, 2021, Kumar et al., 2020, Bastek et al., 2022, Zheng et al., 2021, Mozaffar et al., 2019, Bonatti and Mohr, 2021, Vlassis et al., 2020, Kumar and Kochmann, 2022, Meng et al., 2025]. Several ANN-based constitutive modeling works incorporate novel measures (such as input-convexity) to ensure that the strain-energy density or stress-history predicted by the ANNs are thermodynamically consistent [Huang et al., 2022, Linka and Kuhl, 2023, Masi and Stefanou, 2022], and obey other physical constraints, such as objectivity, material frame indifference and polyconvexity [Klein et al., 2022, Geuken et al., 2025]. Furthermore, to address the lack of availability of full stress-tensor data for training supervised training, several frameworks leverage the balance of linear momentum [Thakolkaran et al., 2022] or the virtual fields method [Meng et al., 2025, Lourenço et al., 2024] to train ANNs or Recurrent Neural Networks (RNNs) to surrogate the stress-strain relationship for the materials using only experimentally available surface displacement and boundary-averaged force data.
@@
Although stress-unsupervised ANN/RNN surrogates are capable of learning complicated constitutive models, they do not provide interpretable constitutive models. The EUCLID framework [Flaschel et al., 2021a] was initially introduced to model complicated homogeneous hyperelastic constitutive behavior in an interpretable manner. This was achieved through sparse symbolic regression using a pre-defined library of known strain energy densities. The EUCLID framework has since been extended to model homogeneous elastoplastic materials [Flaschel et al., 2022], viscoelastic materials [Marino et al., 2023], and generalized standard materials [Flaschel et al., 2023]. Most of the aforementioned works focus on modeling homogeneous materials.
@@
An earlier work, which used only surface displacement data [Guo et al., 2020], focused on elastography of incompressible linear elastic heterogeneous solids, utilizing novel Physics-Informed Neural Networks (PINNs) to predict the variation of Young’s Modulus as a function of position within the material. A later work by the same authors employed multiple PINNs to obtain the spatial variation of Poisson’s ratio in elastography [Chen and Gu, 2023]. Another work presented a detailed study on segmenting heterogeneous hyperelastic materials using surface displacement data and associated strain invariants [Nguyen and Lejeune, 2024a]. A more recent work used a stress-unsupervised ANN-based framework [Shi et al., 2025] to segment and model a heterogeneous hyperelastic solid, wherein the inputs to the ANN are the strain invariants and the query position, and the output of the ANN is the strain energy density. The novel ANN-based framework was able to segment and reproduce the constitutive models of each segment in the hyperelastic solid. However, the use of ANNs makes the discovered models uninterpretable and provides no uncertainty estimates on the discovered models. Furthermore, since the query position is an essential input to the aforementioned ANN, it would complicate the adoption of the ANN as a constitutive model for forward simulation on other geometries. In this context, we propose the Hetero-EUCLID framework, wherein we extend the Bayesian-EUCLID framework [Joshi et al., 2022] to segment a heterogeneous solid and provide interpretable constitutive models for all constituents, using only surface displacement and boundary-averaged force data.
@@
The problem of constitutively modeling all the constituents in a heterogeneous geometry can be divided into two subtasks: segmenting the geometry and obtaining the constitutive model for all the materials involved. In section 2 of the current work we present
@@
  • An overview of extending the Bayesian-EUCLID framework [Joshi et al., 2022], which was originally developed for homogeneous materials, to model segmented homogeneous materials.
@@
A schematic illustration of the proposed Hetero-EUCLID framework is provided in Figure 1. In section 3, we present results of validation studies on square geometries containing two or more hyperelastic constituents. The validation studies utilize surface displacement data and boundary-averaged normal reaction force data from biaxial tension tests on thin hyperelastic specimens.
Figure 1: Schematic of the Hetero-EUCLID workflow for unsupervised constitutive model discovery in heterogeneous hyperelastic materials. (a) illustrates the specimen subjected to non-equi-biaxial tensile loading. (b) and (c) show the point-wise displacement observations and corresponding finite element interpolation onto a wedge-type mesh, respectively. (d) depicts the resulting full-field displacement field, while (e) presents the corresponding strain field obtained via spatial differentiation. (f) represents a catalogue of the hyperelastic constitutive models library. (g) outlines the homogenized model formulation incorporating the weak form of the linear momentum balance, which is solved using ordinary least squares regression to yield the residual force norm distribution shown in (h). Material segmentation based on the residual force indicator is performed using a random seed-based island growth algorithm, as illustrated in (i). (j) presents the unified global heterogenized system of equations assembled after material segmentation. The unknown constitutive model coefficients $\boldsymbol{\theta}^{(i)}$ are solved for using the Bayesian EUCLID framework. (k) highlights the discovered region-wise interpretable constitutive models corresponding to each material segment within the heterogeneous specimen.
For the scope of this study, we source displacement and boundary-averaged normal force data from three-dimensional finite element simulations. This is in contrast to previous similar studies, which focused on validating their results in 2D plane strain settings [Thakolkaran et al., 2022, Shi et al., 2025]. In section 3, we also validate the proposed Hetero-EUCLID framework on “Non-native” meshed geometries, wherein the Finite Element mesh used to perform the inverse Hetero-EUCLID analysis was different from the Finite Element mesh used to generate the displacement and boundary-averaged force. The paper concludes in section 4 with comments on the results of the study and potential future work.
@@ 2.1 Incorporating heterogeneity
Figure 2: Meshing the given thin plate geometry with C3D6 elements with thickness identical to that of the plate.
@@
where $\nabla(\cdot)=\partial(\cdot)/\partial\boldsymbol{X}$ is defined as the gradient operator in the reference coordinates. Due to the use of linear shape functions in the wedge mesh type, $\nabla\mathbb{N}^{a}(X)$ remains constant within each element. For the scope of this work, a single-element-thick mesh is generated over the reference domain. This would imply that all nodes are located on the surface and their displacements $\boldsymbol{u}^{a,t}$ could be obtained through 3D-Digital Image Correlation (DIC) experiments [Luo et al., 1993]. Although single-surface 3D-DIC experiments are somewhat routinely performed, it is challenging to simultaneously perform 3D-DIC to obtain the 3D displacement field on both– the top and bottom surfaces of a thin deforming sample. Such dual-surface displacement field acquisition can be performed using dual-camera stereo-DIC, dual-surface DIC, or mirror-assisted multi-view DIC systems [Takada et al., 2023, Genovese et al., 2021, Sun et al., 2023, Chen and Pan, 2022]. Furthermore, for brevity and without loss of generality, we employ only the data corresponding to a single large deformation state, $t=1$. We will henceforth be dropping the use of $t$ in the superscript.
@@
We now consider a reference domain composed of multiple hyperelastic material segments. Since each segment belongs to the hyperelastic material class, they can be assigned a strain energy density $W(\boldsymbol{F})$ and the corresponding first Piola–Kirchhoff stress tensor would be $\boldsymbol{P}(\boldsymbol{F})=\partial W/\partial\boldsymbol{F}$. In order to satisfy the material objectivity constraint, the strain energy density $W(\boldsymbol{F})$, will be a function of the right Cauchy-Green deformation tensor $\boldsymbol{C}=\boldsymbol{F}^{T}\boldsymbol{F}$ and can be defined as
$W(\boldsymbol{C})=W(I_{1},I_{2},I_{3},I_{a})$ (3)
$I_{1}(\boldsymbol{C})=\mathrm{tr}(\boldsymbol{C}),\quad I_{2}(\boldsymbol{C})=\frac{1}{2}[\mathrm{tr}^{2}(\boldsymbol{C})-\mathrm{tr}(\boldsymbol{C}^{2})],\quad I_{3}(\boldsymbol{C})=\mathrm{det}(\boldsymbol{C}),\quad I_{a}(\boldsymbol{C},\boldsymbol{a})=\boldsymbol{a}\cdot\boldsymbol{Ca}$ (4)
where $I_{1},I_{2},I_{3}$ are isotropic invariants of right Cauchy-Green deformation tensor $\boldsymbol{C}$ and $I_{a}$ represents anisotropic invariant of $\boldsymbol{C}$ with fiber orientation angle $\alpha$ with $\boldsymbol{a}=(\cos{\alpha},\sin{\alpha},0)^{T}$.
@@
We address the inverse problem of identifying the strain energy density function for a specimen subjected to a non-equi-biaxial loading test by introducing a material feature library $\boldsymbol{Q}:\mathbb{R}^{4}\rightarrow\mathbb{R}^{\mathit{n_{f}}}$, where $\mathit{n_{f}}$ denotes the number of nonlinear feature functions. Accordingly, the constitutive relation in Equation 3 can be reformulated using this feature library $\boldsymbol{Q}$ and a vector of unknown material parameters $\boldsymbol{\theta}\in\mathbb{R}^{\mathit{n_{f}}}_{+}$, consisting of non-negative scalar coefficients [Joshi et al., 2022].
$W(I_{1},I_{2},I_{3},I_{a};\boldsymbol{\theta})=\boldsymbol{Q}^{T}(I_{1},I_{2},I_{3},I_{a})\>\boldsymbol{\theta}$ (5)
@@
For the scope of this work, we choose a feature library which can be used to represent the compressible hyperelastic material models such as the Neo-Hookean [Treloar, 1943], Isihara [Isihara et al., 1951], Biderman [Biderman, 1958], Haines-Wilson [Haines and Wilson, 1979], Ogden [Ogden, 1972] and Holzapfel-Gasser-Ogden Holzapfel et al. [2000] models:
@@
$\boldsymbol{Q}\left(I_{1},I_{2},I_{3},I_{a}\right)$ $=\underbrace{\left[\left(\tilde{I}_{1}-3\right)^{i}\left(\tilde{I}_{2}-3\right)^{j-i}:j\in\{1,\dots,N_{\text{poly}}\},i\in\{0,\dots,j\}\right]^{T}}_{\text{Generalized Mooney-Rivlin features }}$
@@
$\quad\oplus\quad\underbrace{\left[(J-1)^{2k}:k\in\{1,\dots,N_{\text{vol}}\}\right]^{T}}_{\text{Volumetric energy features }}\quad\oplus\quad\underbrace{\left[(\tilde{I}_{a}-1)^{l}:l\in\{2,\dots,N_{\text{aniso}}\}\right]^{T}}_{\text{Anisotropy features }}$ (6)
where $J=\mathrm{det}(\boldsymbol{F})=I_{3}^{1/2}$ and $\tilde{I}_{1}=J^{-2/3}I_{1},\>\tilde{I}_{2}=J^{-4/3}I_{2}$ are the isotropic deviatoric invariants of the unimodular Cauchy-Green deformation tensor $(\boldsymbol{\tilde{C}}=J^{-2/3}\boldsymbol{C})$, $\tilde{I}_{a}=J^{-2/3}(\boldsymbol{a}\cdot\boldsymbol{Ca})$ is the anisotropic deviatoric invariant of $\boldsymbol{\tilde{C}}$ and the symbol $\oplus$ represents the concatenation of the features of the chosen library. The physical constraints such as objectivity, material symmetry, and a stress-free reference configuration ($\boldsymbol{P}(\boldsymbol{F}=\boldsymbol{I})=0$) are automatically satisfied [Joshi et al., 2022] by Equation 5 and 6. It should be noted that any terms involving $(\tilde{I}_{2}-3)$ in the feature library $\boldsymbol{Q}$ are not polyconvex, and should ideally be replaced with $(\tilde{I}_{2}^{3/2}-3^{3/2})$ [Hartmann and Neff, 2003]. However, we choose to retain the more traditional $(\tilde{I}_{2}-3)$ term, which appears in the Isihara model [Isihara et al., 1951]. The addition of this term to the feature library makes it convenient to benchmark the Hetero-EUCLID framework using synthetic surface displacement data from commercial finite element software that utilizes the Isihara model. If the $(\tilde{I}_{2}^{3/2}-3^{3/2})$ term were to be included in the feature library instead of the $(\tilde{I}_{2}-3)$ term, then the non-negativity of the entries of $\boldsymbol{\theta}$ automatically ensures that the strain energy density $W(\boldsymbol{F})$ will be polyconvex.
@@
To compute the partial derivatives of the various features in the material library, $\partial\boldsymbol{Q}^{T}/\partial F_{i\mathrm{J}}$, we use the automatic differentiation [Baydin et al., 2018] feature in the torch library of Python. An advantage of using the weak form of the linear momentum balance is that it inherently reduces the sensitivity issues associated with the second-order spatial derivatives present in the strong form [Flaschel et al., 2021b].
@@
Now, we choose various nodal values of the test function $\boldsymbol{v}^{a}$ to generate equations to solve for the constitutive properties of the material
@@
    Within each integral, the material model $\boldsymbol{\theta}\left(\boldsymbol{X}\right)=\boldsymbol{\theta^{(k)}},\text{for }\boldsymbol{X}\in\Omega^{(k)};k\in\{1,\dots,n_{c}\}$. Since $\boldsymbol{\theta^{(k)}}$ is constant in each material segment, it can be taken out of the integral. Equation 13 can thus be rewritten as
@@
    $(v_{i}^{a})_{k}=\begin{cases}1&\forall\>(a,i)\in\mathcal{D}^{\mathrm{fix}}_{k}\\ 0&\text{otherwise}\end{cases}\quad\text{for }k=\{1,\dots,n_{b}\}$
    where, $\mathcal{D}_{k}^{\text{fix}}\subset\mathcal{D}^{\text{fix}}$ denotes the set of degree of freedoms belonging to the $k^{\text{th}}$ boundary $\partial\Omega_{k}$ and $n_{b}$ denotes the total number of boundaries where force is measured. Substituting the value of chosen test function $(v_{i}^{a})_{k}$, Equation 10 gets modified and results in $n_{b}$ set of equations (one at every boundary) as follows
@@
    Using Equation 8 and the methodology adopted in Equation 1315, Equation 17 can be written as